OpenStudy (anonymous):

Use the quadratic formula to solve the equation. –4x^2 – 3x + 2 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (jiteshmeghwal9):

Here, a=-4 b=-3 c=2 therefore,\[\LARGE{-(-3) \pm \sqrt{(-3)^2-4(-4)(2)}\over2(-4)}\]

OpenStudy (jiteshmeghwal9):

solve this :)

OpenStudy (anonymous):

Adding on to hartnn's equation, a=-4, b=-3 & c=2 .

OpenStudy (anonymous):

So just plug and solve

OpenStudy (anonymous):

Adding on to hartnn's equation, a=-4, b=-3 & c=2 .

OpenStudy (anonymous):

ohh , i got it . its the 1st one. thanks yall ! (;

OpenStudy (jiteshmeghwal9):

yw :)